OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation. -4(x + 3) = -4x - 12 A. Commutative Property of Addition B. Associative Property of Addition C. Distributive Property D. Associative Property of Multiplication

OpenStudy (anonymous):

Commutative Prop. of Addition: (order) 3+4=4+3 Associative Prop. of Addition: (grouping) 4+(5+6)=(4+5)+6 Distributive Property: 3(x+4)=3x+12 Associative Prop. of Multiplication: (grouping) 4*(5*6)=(4*5)*6

Directrix (directrix):

Here's a property review chart.

OpenStudy (anonymous):

so Distributive Property

Directrix (directrix):

Yes.

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

No problem!